An element in its ground state has electron configuration: [Ar]4s23d104 What is this element? Answer: Check
Answer #1

Answer:

Ground State Electronic Configuration : [Ar] 4s2 3d10 4p4.

Ar is symbol for Argon nobel gas with 18 hence

[Ar] \equiv 18 electrons and

\therefore Total Number of electrons = 18 + 2 "4s electrons"+ 10 "3d electrons"+ 4 "4p electrons".

\therefore Total Number of electrons = 18 + 2 + 10 + 4 = 34.

With 34 electrons its a Selenium element.

The element is SELENIUM (Se)

====================XXXXXXXXXX=================